# ADPO: Anchored Direct Preference Optimization A Unified Framework from Pairwise to Listwise Preferences

### v5

@@ 3.4 Formal Statements of Core Results
The anchor $\pi_{\mathrm{ref}}$ defines the coordinate origin, while the target $q$ determines the minimizer $u^{\star}$ and local curvature. This provides a distribution-space analogue of TRPO but with an implicit (Fisher-metric) trust region—no explicit KL constraint needed.
Figure 1: ADPO’s anchored trust region. The Fisher metric $\operatorname{Diag}(q)-qq^{\top}$ induces a quadratic “bowl” in the distribution space, expressed in anchored coordinates $u=(s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}$; the quadratic well is centered at $u^{\star}$ (i.e., at the target $q$), not at the anchor itself. The green trajectory shows the optimization path descending toward the optimal policy, with the anchored coordinate system providing an implicit trust region that prevents large distributional drift. This geometric structure holds for both fixed anchors (stable centers for distillation) and dynamic anchors (adaptive regions for exploration).
3.5 KL-Mirror Descent Interpretation
Theorem 3.7 (KL-mirror descent / Bregman-prox step).
If one directly optimizes $\mathcal{L}(u)=A(u)-\langle q,u\rangle$ over anchored logits $u$ (where $A(u)=\log\sum_{j}e^{u_{j}}$), then for any step size $\eta>0$, the one-step update is equivalent to:
$u_{t+1}\ =\ \arg\min_{u}\ \Big\langle\nabla\mathcal{L}(u_{t}),\ u\Big\rangle\ +\ \frac{1}{\eta}\,\mathrm{KL}\!\Big(\tilde{p}(u_{t})\,\big\|\,\tilde{p}(u)\Big),$ (13)
whose solution is the multiplicative-weights update $\tilde{p}_{t+1}\propto\tilde{p}_{t}\odot\exp(-\eta(\tilde{p}_{t}-q))$.
Proof.
A mirror descent step with generator $A$ minimizes $\langle g_{t},u\rangle+\tfrac{1}{\eta}D_{A}(u\|u_{t})$. Using the Bregman divergence relation $D_{A}(u\|u_{t})=\mathrm{KL}(\tilde{p}(u_{t})\|\tilde{p}(u))$ for the softmax family yields (13). ∎∎
3.6 Dynamic Anchor Stability
The Fisher metric $\operatorname{Diag}(q)-qq^{\top}$ is anchor-agnostic (depends only on $q$), allowing flexible anchor strategies: (i) fixed $\pi_{\text{ref}}$ (frozen, stable origin); (ii) dynamic $\pi_{\theta_{\text{old}}}$ (periodic updates, adaptive like TRPO); (iii) teacher $\pi_{\text{teacher}}$ (offline distillation). We characterize monotonic improvement within each fixed-anchor phase:
Theorem 3.8 (Fixed-anchor-phase monotonic descent).
@@ 3.6 Dynamic Anchor Stability
When the anchor is fixed, $\mathcal{L}(u)=A(u)-\langle q,u\rangle$ is $\mu$-strongly convex on the quotient space and $L$-smooth on any bounded domain where $\tilde{p}_{\theta}(u)_{i}\in[\varepsilon,1-\varepsilon]$. Consequently, gradient descent with step size $\eta\in(0,1/L]$ (or KL-mirror descent with the corresponding step-size condition) enjoys the linear rate
$\mathcal{L}(u_{t+1})-\mathcal{L}(u^{\star})\leq(1-\mu\eta)\big(\mathcal{L}(u_{t})-\mathcal{L}(u^{\star})\big)$
within that fixed-anchor phase. When the anchor updates ($\pi_{\mathrm{ref}}\leftarrow\pi_{\theta}$), the coordinate system changes—we do not claim monotonicity across anchor updates.
@@
$A(u)$ is strictly convex with Hessian $\nabla^{2}A(u)=\operatorname{Diag}(\tilde{p}_{\theta}(u))-\tilde{p}_{\theta}(u)\tilde{p}_{\theta}(u)^{\top}$ (softmax Fisher metric). Near $u^{\star}$ where $\tilde{p}_{\theta}(u^{\star})=q$, this coincides with $\operatorname{Diag}(q)-qq^{\top}$ up to higher-order terms, yielding local strong convexity. In bounded domains where $\tilde{p}_{\theta}(u)_{i}\in[\varepsilon,1-\varepsilon]$ for all $i$, the Fisher metric satisfies $\nabla^{2}\mathcal{L}(u)\succeq\varepsilon\,P_{\mathbf{1}^{\perp}}$ globally in this domain (one may take $c(\varepsilon)=\varepsilon$ in Corollary 3.6). The spectral upper bound of $\nabla^{2}A(u)$ is also bounded in this domain, yielding $L$-smoothness with $L=L(\varepsilon,K)$. ∎∎
Implication. Within each fixed-anchor phase, Fisher-metric trust region ensures monotonic descent; across anchor updates, the coordinate system adapts (like TRPO). Experiments (Section 5) reveal task-dependent trade-offs: dynamic anchors excel at online exploration; fixed anchors excel at offline distillation.
Gradients.
Pairwise: $\frac{\partial\ell_{ij}}{\partial\theta}=\beta\big(\sigma(\beta(\Delta_{\theta}-\Delta_{\mathrm{ref}}))-q_{ij}\big)\big(\nabla_{\theta}s_{i}-\nabla_{\theta}s_{j}\big).$
Listwise: $\frac{\partial\mathcal{L}_{\mathrm{group}}^{\mathrm{ref}}}{\partial s_{i}}=\frac{1}{\tau_{\mathrm{anc}}}\big(\tilde{p}_{\theta}(i\mid S_{x})-q(i\mid S_{x})\big).$
@@ 4 Experimental Setup
  • WinMass: For each context $x$ with candidate set $S_{x}=\{1,\ldots,P\}$ and ground-truth ranking (higher reward is better), WinMass measures the total probability mass assigned to the top item under the anchored distribution: $\mathrm{WinMass}=\mathbb{E}_{x}[\tilde{p}_{\theta}(\operatorname*{arg\,max}_{i}R_{i}^{\star}\mid S_{x})]$. Higher is better; random baseline = $1/P$.
  • NDCG@k: Normalized Discounted Cumulative Gain at position $k$, a standard ranking metric that rewards placing high-reward items at top positions with logarithmic discounting. Higher is better; range $[0,1]$.
  • KL divergence: For distillation tasks, we measure $\mathrm{KL}(\pi_{\mathrm{teacher}}\|\pi_{\mathrm{student}})$, the expected KL divergence from teacher to student policy over states and actions. Lower is better, indicating closer alignment with the teacher.
@@
The $2\times 2$ base covers pairwise methods (4 combinations), plus 3 ADPO listwise extensions (Raw/KDE/KDE-Rank aggregating full distributions), yielding 7 methods total: Standard DPO Pairwise-Soft/Hard, ADPO Pairwise-Soft/Hard, ADPO Listwise-Raw/KDE/KDE-Rank.
@@
    Heavy Noise: Gaussian noise with outliers. Light (SNR=1.0, 5% outliers), Medium (SNR=0.5, 10%), Heavy (SNR=0.2, 20%).
@@
Model architecture. Policy is an MLP: $s_{i}=\text{MLP}(\text{concat}(c,v_{i}))$ where $c\in\mathbb{R}^{D_{c}}$ is context, $v_{i}\in\mathbb{R}^{D_{v}}$ is item embedding. We test 3 scales:
@@ 5.1 $2\times 2$ Base Comparison Across Scenarios
Across all 12 scenarios, ADPO shows consistent relative gains over the standard DPO baseline, with improvements ranging from 12% to 79% (10 seeds). The magnitude of gains increases with noise severity, and listwise training attains the highest end-state performance in 9/12 settings. Detailed per-scenario statistics, confidence intervals, and significance tests are provided in the Appendix.
Table 3: WinMass across 12 scenarios (random baseline = 1/P = 0.25 for P=4). Values are mean over 10 random seeds (std << 0.05 for most entries; 95% CI and Wilcoxon p-values in Appendix). Bold: best method. Underline: best among pairwise. "Best Listwise" shows the highest-performing variant among ADPO Listwise-Raw/KDE/KDE-Rank, with superscript: $^{\text{R}}$=Raw, $^{\text{K}}$=KDE, $^{\text{KR}}$=KDE-Rank. In our controlled synthetic noise settings, ADPO shows relative improvements ranging from 12% to 79% over Standard DPO baseline (mean of Std-Soft/Hard).
@@
Figure 2: Comprehensive $2\times 2$ comparison across 12 scenarios (10 seeds each). Each subplot shows convergence curves for 7 methods. Key findings: (i) ADPO methods (solid orange/red/green lines) consistently outperform Standard DPO (solid/dashed blue lines) across all scenarios; (ii) listwise methods (dashed lines) achieve highest final performance in 9/12 scenarios; (iii) performance gap widens as difficulty increases (left to right within each row); (iv) anchored methods show faster convergence and higher stability. Error bands: mean $\pm$ s.e.
@@
Listwise methods (ADPO Listwise-Raw/KDE/KDE-Rank) achieve the highest WinMass in 9 out of 12 scenarios (Table 3). Peak performance: 0.849 (Distribution Shift-Medium, Listwise-Raw). Listwise methods:
@@
Table 4: Soft vs. hard label comparison (pairwise methods only). Winner highlighted. Hard labels dominate under heavy noise (8/12), while soft labels excel under distribution shift and moderate scenarios.
@@
Heavy Noise Heavy 0.430 0.431 0.702 0.770
Dist. Shift Light 0.712 0.740 0.841 0.794
Adversarial Medium 0.629 0.658 0.748 0.730
Heavy-Tailed Heavy 0.458 0.472 0.784 0.746
Hard wins: 8/12 scenarios (Heavy Noise all, Adversarial 2/3, Dist. Shift 2/3)
Soft wins: 4/12 scenarios (Heavy-Tailed 2/3, Adversarial 1/3, Dist. Shift 1/3)
Contrary to conventional wisdom, hard labels dominate under heavy noise (Table 4). Under Heavy Noise-Heavy, ADPO-Hard achieves 0.770 vs. ADPO-Soft’s 0.702 (+9.7%). However, soft labels excel under distribution shift (Light: 0.841 vs. 0.794, +5.9%) and moderate adversarial scenarios. Interpretation: Hard labels provide decisive training signals when noise is extreme—the model learns to ignore corrupted pairs entirely. Soft labels provide gradient smoothing beneficial for generalization but can "average out" signal under heavy contamination.

### v6

@@ 3.5 Anchor Strategies: Offline vs Online
Key property: The anchoring defines a moving coordinate system, while the Fisher metric (determined by $q$) provides an implicit trust region. This mirrors TRPO’s $\mathrm{KL}(\pi_{\text{old}}\|\pi_{\text{new}})$ constraint [23] but operates in distribution space with groupwise shift invariance.
Task-dependent trade-off.
Our experiments (Section 5) reveal:
  • Dynamic anchors improve online exploration in noisy environments by 5–11% in 3/4 scenarios (Section 5.4). The adaptive coordinate frame enables effective exploration while maintaining stability.
[…]
Remark 3.9 (Connection to existing methods).
Offline: No anchor recovers SFT/KD; fixed anchor adds regularization toward a reference policy. Online: Dynamic anchor connects to TRPO/PPO but with implicit Fisher-metric trust regions and groupwise shift invariance.
3.6 Stability and Convergence Analysis
Having established the geometric structure (Section 3.3) and algorithmic view (Section 3.4) of ADPO, we now analyze its stability and convergence properties.
Fixed anchor: Strong convergence guarantees.
When the anchor $\pi_{\mathrm{ref}}$ is fixed throughout training, ADPO enjoys strong convexity and linear convergence within the anchored coordinate system.
Theorem 3.10 (Monotonic descent under fixed anchor).
@@ 3.6 Stability and Convergence Analysis
With a fixed anchor, $\mathcal{L}(u)=A(u)-\langle q,u\rangle$ is $\mu$-strongly convex and $L$-smooth on the bounded domain where $\tilde{p}(u)\in[\varepsilon,1-\varepsilon]^{P}$. For step size $\eta\in(0,1/L]$, gradient or mirror descent satisfies
$\mathcal{L}(u_{t+1})-\mathcal{L}(u^{\star})\;\leq\;(1-\mu\eta)\big(\mathcal{L}(u_{t})-\mathcal{L}(u^{\star})\big),$ (18)
yielding linear convergence within each fixed-anchor phase.
@@
The Hessian $\nabla^{2}A(u)=\operatorname{Diag}(\tilde{p})-\tilde{p}\,\tilde{p}^{\top}$ has eigenvalues in $\min_{i}\tilde{p}_{i}(1-\tilde{p}_{i}),\max_{i}\tilde{p}_{i}(1-\tilde{p}_{i})]$. On the bounded domain $\varepsilon,1-\varepsilon]^{P}$, this yields:
  • Strong convexity: $\mu\geq\varepsilon(1-\varepsilon)$
  • Smoothness: $L\leq 1/4$
Standard convergence rates for strongly convex and smooth functions then give the stated linear convergence [28]. ∎
Corollary 3.11 (Convergence rate).
With optimal step size $\eta=1/L$, the number of iterations to reach $\varepsilon$-accuracy is:
$T=O\!\Big(\frac{L}{\mu}\log\frac{1}{\varepsilon}\Big)\;=\;O\!\Big(\frac{1}{\varepsilon(1-\varepsilon)}\log\frac{1}{\varepsilon}\Big).$
For well-separated distributions (large $\varepsilon$), convergence is fast.
Dynamic anchor: Coordinate frame changes.
When the anchor updates ($\pi_{\mathrm{ref}}\!\leftarrow\!\pi_{\theta}$), the coordinate frame changes and monotonicity is not guaranteed across updates. However, local stability is preserved.
Theorem 3.12 (Local stability under dynamic anchor).
Let $\pi_{\mathrm{ref}}^{(t)}$ denote the anchor at iteration $t$, and define anchored logits $u^{(t)}=(\ell_{\theta}-\ell_{\mathrm{ref}}^{(t)})/\tau_{\mathrm{anc}}$. Within each anchor phase $t_{k},t_{k+1})$, the loss $\mathcal{L}^{(k)}(u^{(t)})$ decreases monotonically (Theorem 3.10).
At anchor update $t_{k+1}$, the coordinate system shifts, and the loss $\mathcal{L}^{(k+1)}(u^{(k+1)})$ is measured in the new coordinate frame. Global monotonicity $\mathcal{L}^{(k+1)}\leq\mathcal{L}^{(k)}$ is not guaranteed, but the Fisher-metric constraint ensures local stability: the policy remains within a trust region around the previous policy.
Remark 3.13 (Comparison to TRPO/PPO).
TRPO [23] enforces monotonic improvement through explicit KL constraints, while PPO [4] uses clipping to approximate this. Online ADPO achieves similar stability through implicit Fisher geometry, without requiring explicit constraints or clipping. The key difference is that ADPO operates in distribution space (anchored coordinates), while TRPO/PPO operate in parameter space.
3.7 Algorithm and Implementation
We now provide a unified algorithmic description of ADPO that covers both online RL and offline preference learning through a mode parameter.
Algorithm: ADPO (Unified).
Algorithm 1 ADPO (Unified): Online RL and Offline Preference Learning
1:Mode $\in\{\textsc{online},\textsc{offline}\}$; temperature $\tau_{\mathrm{anc}}$; learning rate(s) $\eta_{\pi}$ (and $\eta_{Q}$ for online); anchor update period $N_{\mathrm{anchor}}$; reward temperature $\beta_{r}$ (online only)
2:Initialize policy $\pi_{\theta}$; set anchor $\pi_{\mathrm{ref}}\leftarrow\pi_{\theta}$
3:if online then
4:  Initialize critic $Q_{\phi}$, replay buffer $\mathcal{B}$
5:else
6:  Load dataset $\mathcal{D}=\{(x,S_{x},\text{supervision})\}$
7:end if
8:for training loop $t=1,\ldots,T$ $\triangleright$ iteration for online; epoch for offline do
9:  if online then
10:   Collect trajectories using $\pi_{\theta}$, add to $\mathcal{B}$
11:  end if
12:  for batch update do
13:   if online then
14:     Sample states $\{x_{j}\}_{j=1}^{B}\sim\mathcal{B}$
15:   else
16:     Sample batch $\{(x_{j},S_{x_{j}},\text{sup}_{j})\}_{j=1}^{B}\sim\mathcal{D}$
17:   end if
18:   for each context $x_{j}$ in batch do
19:     // Build target distribution $q$
20:     if online then
21:      Sample candidates $\{y_{1},\ldots,y_{P}\}$: $y_{i}\sim\pi_{\theta}(\cdot|x_{j})$ or mixed sampling
22:      Compute scores: $R_{i}\leftarrow Q_{\phi}(x_{j},y_{i})$ or $r(x_{j},y_{i})$
23:      Normalize (Eq. 19): $\tilde{R}_{i}\leftarrow(R_{i}-\mathrm{median}(R))/(\mathrm{MAD}(R)+\varepsilon)$
24:      Build target: $q(i|x_{j})\leftarrow\mathrm{softmax}(\tilde{R}/\beta_{r})_{i}$
25:     else
26:      Use candidate set $S_{x_{j}}$ from data
27:      Compute target: $q(i|S_{x_{j}})$ from Bradley-Terry, Plackett-Luce, reward model, etc.
28:     end if
29:     $\triangleright$ ===== Core ADPO Update =====
30:     Compute log-probabilities: $\ell_{i}\leftarrow\log\pi_{\theta}(y_{i}|x_{j})$, $\ell_{i}^{\mathrm{ref}}\leftarrow\log\pi_{\mathrm{ref}}(y_{i}|x_{j})$
31:     $u_{i}\leftarrow(\ell_{i}-\ell_{i}^{\mathrm{ref}})/\tau_{\mathrm{anc}}$
32:     $q$-centering (Eq. 10): $\bar{u}_{i}\leftarrow u_{i}-\sum_{k\in S_{x_{j}}}q(k|x_{j})\,u_{k}$
33:     $\tilde{p}_{\theta}(i|x_{j})\leftarrow\mathrm{softmax}(\bar{u})_{i}$
34:   end for
35:   Compute loss: $\mathcal{L}_{\pi}\leftarrow-\frac{1}{B}\sum_{j=1}^{B}\sum_{i\in S_{x_{j}}}q(i|x_{j})\log\tilde{p}_{\theta}(i|x_{j})$
36:   Update policy: $\theta\leftarrow\theta-\eta_{\pi}\nabla_{\theta}\mathcal{L}_{\pi}$
37:  end for
38:  if online then
39:   Update critic $Q_{\phi}$ using standard RL (e.g., TD, SAC)
40:  end if
41:  // Update anchor
42:  if $t\bmod N_{\mathrm{anchor}}=0$ then
43:   $\pi_{\mathrm{ref}}\leftarrow\pi_{\theta}$ $\triangleright$ dynamic anchor (TRPO-style)
44:  end if
45:end for
46:return $\pi_{\theta}$
Practical notes.
  • $q$-centering (always): Apply Eq. 10 to center anchored logits at target $q$, ensuring numerical stability and Fisher-metric consistency.
  • Reward normalization (online RL only): In online mode (Algorithm 1), standardize rewards before building $q$:
    $\tilde{R}_{i}=\frac{R_{i}-\mathrm{median}(R)}{\mathrm{MAD}(R)+\varepsilon},\qquad q(i|x)=\mathrm{softmax}(\tilde{R}/\beta_{r}),$ (19)
    where $\mathrm{MAD}(R)=\mathrm{median}(|R_{i}-\mathrm{median}(R)|)$. Not needed in offline mode where $q$ is fixed.
  • Anchor caching: Cache $\ell^{\mathrm{ref}}$ on batch data to avoid redundant forward passes.
  • Numerical stability: Use log-sum-exp trick for $\log\sum_{j}\exp(u_{j})$.
  • Temperature: In KD scenarios, $\tau_{\mathrm{anc}}\equiv T$ (teacher temperature); see Table 1.
  • Batch processing: Process groups independently; no cross-group dependencies.
@@ 4 Experimental Setup
  • WinMass: For each context $x$ with candidate set $S_{x}=\{1,\ldots,P\}$ and ground-truth ranking (higher reward is better), WinMass measures the total probability mass assigned to the top item under the anchored distribution: $\mathrm{WinMass}=\mathbb{E}_{x}[\tilde{p}_{\theta}(\operatorname*{arg\,max}_{i}R_{i}^{\star}\mid S_{x})]$. Higher is better; random baseline = $1/P$. Computed over 1000 test contexts per scenario.
  • NDCG@k: Normalized Discounted Cumulative Gain at position $k$, a standard ranking metric that rewards placing high-reward items at top positions with logarithmic discounting. Higher is better; range $[0,1]$. We report NDCG@4 for $P=4$ candidates.
  • KL divergence: For distillation tasks, we measure $\mathrm{KL}(\pi_{\mathrm{teacher}}\|\pi_{\mathrm{student}})$, the expected KL divergence from teacher to student policy over states and actions. Lower is better, indicating closer alignment with the teacher. Computed over 10K states sampled from the replay buffer.
Reporting standards.
All results report mean $\pm$ standard error over 10 random seeds (5 seeds for continuous control). We verify statistical significance using Wilcoxon signed-rank tests (p<0.05) for all claimed improvements; detailed confidence intervals and p-values are provided in the Appendix. Training steps and sample counts are aligned across all methods within each experiment.
Student–teacher KL.
We report the group-averaged divergence
$\mathbb{E}_{x,S_{x}}\Big[\mathrm{KL}\big(\pi_{T}(\cdot\mid S_{x})\|\pi_{\theta}(\cdot\mid S_{x})\big)\Big],$
where both policies are renormalized on the candidate set $S_{x}$ before computing KL.
@@
The $2\times 2$ base covers pairwise methods (4 combinations), plus 2 ADPO listwise extensions (Raw/KDE aggregating full distributions), yielding 6 methods total: Standard DPO Pairwise-Soft/Hard, ADPO Pairwise-Soft/Hard, ADPO Listwise-Raw/KDE.
@@
    Heavy Noise: Gaussian noise with outliers. Light (noise scale=1.0, 5% outliers), Medium (scale=2.0, 10%), Heavy (scale=2.5, 20%).
@@
Model architecture. Policy is an MLP: $\ell_{i}=\text{MLP}(\text{concat}(c,v_{i}))$ where $c\in\mathbb{R}^{D_{c}}$ is context, $v_{i}\in\mathbb{R}^{D_{v}}$ is item embedding. We test 3 scales:
@@ 5.1 $2\times 2$ Base Comparison Across Scenarios
Across all 12 scenarios, ADPO shows consistent relative gains over the standard DPO baseline, with improvements ranging from 12% to 93% (10 seeds). The magnitude of gains increases with noise severity, and listwise training attains the highest end-state performance in 11/12 settings. Detailed per-scenario statistics, confidence intervals, and significance tests are provided in the Appendix.
Table 2: WinMass across 12 scenarios (random baseline = 1/P = 0.25 for P=4). Values are mean over 10 random seeds (std << 0.05 for most entries; 95% CI and Wilcoxon p-values in Appendix). Bold: best method. Underline: best among pairwise. "Best Listwise" shows the highest-performing variant among ADPO Listwise-Raw/KDE, with superscript: $^{\text{R}}$=Raw, $^{\text{K}}$=KDE. In our controlled synthetic noise settings, ADPO shows relative improvements ranging from 12% to 93% over Standard DPO baseline (mean of Std-Soft/Hard).
@@
Figure 2: Comprehensive $2\times 2$ comparison across 12 scenarios (10 seeds each). Each subplot shows convergence curves for 6 methods. Key findings: (i) ADPO methods (solid orange/red/green lines) consistently outperform Standard DPO (solid/dashed blue lines) across all scenarios; (ii) listwise methods achieve highest final performance in 11/12 scenarios; (iii) performance gap widens as difficulty increases (left to right within each row), reaching 93% improvement under Heavy Noise-Heavy; (iv) anchored methods show faster convergence and higher stability. Error bands: mean $\pm$ s.e.
@@
Listwise methods (ADPO Listwise-Raw/KDE) achieve the highest WinMass in 11 out of 12 scenarios (Table 2). Peak performance: 0.826 (Distribution Shift-Medium, Listwise-Raw). Listwise methods:
@@
Table 3: Soft vs. hard label comparison (pairwise methods only). Winner highlighted. Hard labels dominate under heavy noise (7/12), while soft labels excel under distribution shift and moderate scenarios.
@@
Heavy Noise Heavy 0.375 0.418 0.725 0.675
Dist. Shift Light 0.653 0.717 0.752 0.738
Adversarial Medium 0.636 0.643 0.730 0.744
Heavy-Tailed Heavy 0.513 0.508 0.722 0.714
Hard wins: 7/12 scenarios (Heavy Noise 2/3, Adversarial 2/3, Heavy-Tailed 2/3)
Soft wins: 5/12 scenarios (Dist. Shift 2/3, Heavy Noise 1/3, others)
The soft vs. hard label comparison reveals a nuanced picture (Table 3). Hard labels win in 7/12 scenarios, particularly under Heavy Noise (2/3), Adversarial (2/3), and Heavy-Tailed (2/3) settings. However, soft labels excel under distribution shift (2/3 scenarios) and achieve the single highest improvement: ADPO-Soft reaches 0.725 vs. ADPO-Hard’s 0.675 under Heavy Noise-Heavy (+7.4%). Interpretation: The choice depends on noise type—hard labels provide decisive signals when noise is extreme and random, while soft labels preserve gradient information beneficial for systematic distribution shifts and moderate noise.